OpenStudy (anonymous):

if g(x)=2x2-x then find g(a+h)

OpenStudy (saifoo.khan):

x(2x - 1)

OpenStudy (saifoo.khan):

so g = x a = 2x h = -1

OpenStudy (anonymous):

g(x) = 2x^2-x g(a+h) = 2(a+h)²-a-h = 2(a²+h²+2ah)-a-h = 2a² +h² +4ah -a -h.

OpenStudy (anonymous):

is the answer x(2x-1)?